Solve:
log(36,10x-11)+log(36,X)=1/2
i get to this point here: log(36,10x^2-11x)=1/2
but then im lost after
thanks
Angela

Answer by nerdybill(7384)

Solve:
log(36,10x-11)+log(36,X)=1/2
.
i get to this point here:
log(36,10x^2-11x)=1/2 (so far, so good)
.
36^log(36,10x^2-11x) = 36^(1/2)
10x^2-11x = 36^(1/2)
10x^2-11x = 6
.
10x^2-11x = 6
10x^2-11x - 6= 0
10x^2 -15x+4x - 6= 0
(10x^2-15x) + (4x-6) = 0
5x(2x-3) + 2(2x-3) = 0
(2x-3)(5x+2) = 0
x = {3/2, -2/5}
throw out the negative solution (extraneous) leaving:
x = 3/2